OpenStudy (anonymous):

Use a cofunction to write an expression equal to cos\frac{ 2pi }{ 7 }

OpenStudy (anonymous):

\[\cos\frac{ 2\pi }{ 7 } \]

OpenStudy (ikram002p):

\(\large \cos \frac{2\pi}{7}=\sin (\frac{\pi}{2} -\frac{2\pi}{7})\)

OpenStudy (anonymous):

thank you